OpenStudy (sleepyjess):

If 1.5 moles of copper metal reacts with 4.0 moles of silver nitrate, how many moles of silver metal can be formed, and how many moles of the excess reactant will be left over when the reaction is complete? Unbalanced equation: Cu + AgNO3 → Cu(NO3)2 + Ag Be sure to show all of your work.

OpenStudy (sleepyjess):

Here's what I've done, I just want to make sure I'm doing it right: Balanced equation: Cu + 2AgNO3 --> Cu(NO3)2 + 2Ag 1 mole of copper reacts with 2 moles of silver nitrate so 1.5 moles of copper would react with 3 moles of silver nitrate to produce 3 moles of silver. Copper is the limiting reactant There would be 1 mole of silver nitrate remaining.

OpenStudy (sleepyjess):

@confluxepic @Compassionate @Jhannybean @Australopithecus @perl

OpenStudy (anonymous):

Is this a test or quiz question or homework question?

OpenStudy (sleepyjess):

It's a lesson review question.

OpenStudy (jhannybean):

These are the number of moles NEEDED\[\sf 1.5 ~\text{mol Cu} \cdot \frac{2 ~\text{mol AgNO3}}{1~\text{mol Cu}} =~?\]\[\sf 4.0 ~\text{mol AgNO3} \cdot \frac{1 ~\text{mol Cu}}{2~\text{mol AgNO3}}=~? \] Compare to: Number of moles you HAVE:\[1.5~\text{mol Cu}\]\[4.0~\text{mol AgNO3}\]

OpenStudy (jhannybean):

After calculating the number of moles of each that is needed, compare it to the amount of moles you have, it will be evident which one is the LR and which is the XS

OpenStudy (sleepyjess):

Thank you @Jhannybean !

OpenStudy (jhannybean):

No problem :) Hope it makes sense

OpenStudy (sleepyjess):

It does!

OpenStudy (jhannybean):

Awesome!